# The quantum-refractive evolution of polarization states in pulsar emission

### v1

@@ Abstract
Highly magnetized neutron stars have quantum refraction effects on pulsar emission due to the non-linearity of the quantum electrodynamics (QED) action. In this paper, we investigate the evolution of the polarization states under the quantum refraction effects combined with the frequency dependence of pulsar emission; we solve a system of evolution equations of the Stokes vector, where the birefringent vector, in which such effects are encoded, acts on the Stokes vector. At a fixed frequency of emission, depending on the magnitude of the birefringent vector, dominated mostly by the magnetic field strength, the evolution of the Stokes vector largely exhibits three different patterns: (i) monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviours. These features are understood and confirmed by means of approximate analytical solutions to the evolution equations.
@@
pulsars – magnetic fields – photon emission – non-linear electrodynamics – quantum refraction – polarization evolution
pubyear: 2024pagerange: The quantum-refractive evolution of polarization states in pulsar emissionB
@@ 1 Introduction
Highly magnetized neutron stars are interesting astrophysical objects that will provide an astrophysical arena for probing fundamental physics (Raffelt (1996)), in particular, strong field QED (Meszaros (1992); Harding & Lai (2006)). The magnetic fields of neutron stars of $10^{12}-10^{15}\,\mathrm{G}$ are strong enough to give rise to QED effects, such as vacuum polarization and pair production due to the induced electric fields, which have been proposed as an astrophysical laboratory to test strong field QED (see Kim & Kim (2021) and references therein). A great advantage of neutron stars as a laboratory is virtually infinitely long duration and range of the electromagnetic fields (in the scale of Compton time and length), strongly contrasting with the short, pulsed fields achieved with ultra-intense lasers, as well as the strength of the fields (Yoon et al. (2021)).
Strong electromagnetic fields polarize the vacuum, which leads to the Heisenberg-Euler-Schwinger (HES) one-loop effective action (Heisenberg & Euler (1936); Schwinger (1951)). The effective action describes non-linear electrodynamics, which has an imaginary part from the vacuum instability due to production of charged particle pairs via electric fields. The HES action leads to the vacuum multiple refraction indices, which is generally referred to as the quantum refraction effect. A test of strong-field QED is one of the most important problems in physics, and there have been laboratory experiment proposals for it, e.g., PVLAS (Polarizzazione del Vuoto con LAser) project (Ejlli et al. (2020)). On the other hand, highly magnetized neutron stars provide near-critical fields, and these can be used as various celestial laboratories to test strong field QED (Ruffini et al. (2010); Fedotov et al. (2023); Hattori et al. (2023))). Also, there have been space missions proposed to observe the strong-field QED effects, e.g., the enhanced X-ray Timing and Polarimetry (eXTP) (Santangelo et al. (2019)) and the Compton Telescope project (Wadiasingh et al. (2019)).
@@
The HES action is well approximated by the post-Maxwellian action, even up to the strength one order lower than the critical magnetic field $B_{\mathrm{c}}=m_{\mathrm{e}}^{2}c^{3}/e\hbar=4.4\times 10^{13}\,\mathrm{G}$, which keeps up to the quadratic terms of the Maxwell scalar and pseudo-scalar. Therefore, the post-Maxwellian action exhibits non-linear characteristics of vacuum polarization, such as quantum refraction. Previously, we have studied the quantum refraction effects on the polarization vectors in the dipole magnetic field background of a pulsar model (Kim et al. (2024)). The study is non-trivial in comparison with other similar studies where the background magnetic field is assumed to be uniform, in that we have to deal with dipole magnetic field, whose strength and direction vary over space.
@@
In this paper, we study the evolution of the polarization states under the quantum refraction effects combined with the frequency dependence of pulsar emission. Our analysis is based on the evolution equations of the Stokes vector, where such effects are encoded into the birefringent vector that acts on the Stokes vector. The Stokes vector has a crucial advantage over the polarization vector in representing polarization states: it can be directly determined from experimentally measurable quantities and accommodate depolarization effects due to incomplete coherence and random processes during the photon propagation. Solutions of the evolution equations describe how the polarization states change along the photon propagation path from the emission point towards an observer. It turns out that the evolution of the Stokes vector, at a fixed frequency of emission, largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, dominated mostly by the magnetic field strength: (i) fractionally oscillatory - monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviours, which are found by numerical solutions and also confirmed by approximate analytical solutions. These are novel features rarely illuminated in previous studies on the same topic.
@@
The paper is organized as follows. In Section 2.1, we introduce a system of evolution equations of the Stokes vector and apply this formalism to our pulsar emission model for an oblique rotator. In Section 2.2, the evolution equations are solved for some known rotation-powered pulsars (RPPs) in three ways: fully numerically, via perturbation analysis, and using an analytical approximation. Also, we discuss the evolution patterns of the Stokes vectors resulting from the solutions. Then finally, we conclude the paper with discussions on other similar studies and follow-up studies.
@@
2 Evolution of Stokes vector in strong magnetic field
@@ 2.1 Evolution equations of Stokes vector
Classically, polarization properties of pulsar emission are described by the Stokes parameters $\left\{I,Q,U,V\right\}$, where $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization of the curvature emission in pulsars (for more details, see Appendix A). However, in the presence of strong magnetic fields in the background of the emission, the polarization evolves along the photon propagation path from the emission point towards an observer. The evolution of the polarization can be investigated systematically using the formalism initiated by Kubo & Nagata (1981, 1983, 1985), and further developed by Heyl & Shaviv (2000), Heyl et al. (2003), Heyl & Caiazzo (2018), and Novak et al. (2018), namely, a system of evolution equations of the Stokes vector, described as
@@
where $k\equiv\omega/c$ denotes the wave number for the electromagnetic radiation and $s$ is an affine parameter to measure the length of the photon trajectory, and $\mathbf{S}$ is the normalized Stokes vector, defined out of the Stokes parameters as $\mathbf{S}=\left(S_{1},S_{2},S_{3}\right)\allowbreak\equiv\left(Q/I,U/I,V/I\right)$, [^1] The classical Stokes vector can be expressed via pulse profiles of pulsar curvature emission, as illustrated in Appendix A. and $\mathbf{\hat{\Omega}}$ is the birefringent vector, defined as [^2] Note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in the references above.
@@
with $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ being the two classical mode polarization vectors, orthogonal to each other and to $\mathbf{\hat{n}}_{\left[0\right]}$; the specific forms of $\mathbf{\hat{n}}_{\left[0\right]}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ are later given by equations (8), (12) and (13), respectively, for the magnetic field of an oblique dipole rotator as described by equation (5) and Fig. 1.
Figure 1: A cross-sectional view of a pulsar magnetosphere with the dipole magnetic field lines (green) around a neutron star. The vertical dashed line (black) and the inclined solid line (red) represent the rotation axis and the magnetic axis, respectively. $\alpha$ between these axes denotes the inclination angle. The scale of the unity in this graph is equivalent to the neutron star radius $\sim 10^{6}\,\mathrm{cm}$. The red dotted line represents the trajectory curve of the light ray traced by $\mathbf{\hat{n}}_{\left[0\right]}$ as projected onto the $xz$-plane. (Credit: Kim & Trippe (2021), reproduced with modifications.)
@@
$\mathbf{B}\left(r,\theta,\phi\right)=\frac{2\mu\left(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{r}}+\frac{\mu\left(\cos\alpha\sin\theta-\sin\alpha\cos\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{\theta}}+\frac{\mu\sin\alpha\sin\phi\,}{r^{3}}\mathbf{e}_{\hat{\phi}},$ (5)
@@
However, at the same time, our pulsar magnetosphere rotates, and therefore the field lines get twisted due to the magneto-centrifugal acceleration on the plasma particles moving along the field lines (Blandford & Payne (1982)). Then, taking into consideration this magneto-hydrodynamic (MHD) effect, the direction of the classical photon propagation, which must line up with the particle velocity in order for an observer to receive the radiation, can be described as (Gangadhara (2005))
$\mathbf{\hat{n}}_{\left[0\right]}=\beta\mathbf{\hat{B}}+\mathbf{\Omega}_{\mathrm{p}}\times\mathbf{r,}$ (6)
@@
During the rotation the azimuthal phase changes by $\phi\sim\Omega_{\mathrm{p}}t$, while our photon has propagated a distance by $s\sim ct$. In our analysis, the photon propagation is described with the consideration of the MHD effect above, assuming $\phi$ to be very small; e.g., $\phi\lesssim 10^{-1}$ is considered for a millisecond pulsar with $\Omega_{\mathrm{p}}\sim 10^{2}\,\mathrm{Hz}$, during the time of rotation $t\lesssim 10^{-3}\,\mathrm{s}$, such that $s\lesssim 10^{7}\,\mathrm{cm}$, which corresponds to the propagation distance within about $10$ times the neutron star radius. For equation (6) we take only the leading order expansions of $\mathbf{\hat{B}}\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ and $\beta\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ in $\phi$ from equations (5) and (7), respectively, and can express the classical propagation vector $\mathbf{\hat{n}}_{\left[0\right]}$ in Cartesian coordinates as
@@
where we have considered $\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\lesssim\phi$, e.g., for a millisecond pulsar with $\Omega_{\mathrm{p}}\sim 10^{2}\,\mathrm{Hz}$ and $r_{\mathrm{o}}\sim 10^{6}\,\mathrm{cm}$, such that $\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\lesssim\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\lesssim\phi^{2}$, all to be ignored in our analysis, and have substituted $\phi=\Omega_{\mathrm{p}}s/c$ in equation (11), the leading order rotational effect to be considered in our analysis.
The orthogonal pair of classical mode polarization vectors, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$, both being also orthogonal to $\mathbf{\hat{n}}_{\left[0\right]}$ as given by equation (8) above, are determined as
@@
such that the three vectors, $\mathbf{\hat{n}}_{\left[0\right]}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ form an orthonormal basis. [^5] It can be checked out that $\mathbf{\hat{n}}_{\left[0\right]}\cdot\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}\approx 0+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$, $\mathbf{\hat{n}}_{\left[0\right]}\cdot\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}=0$ and $\boldsymbol{\varepsilon}_{\,\mathrm{I}\left[0\right]}\cdot\boldsymbol{\ \varepsilon}_{\mathrm{II}\left[0\right]}=0$ while $\mathbf{\hat{n}}_{\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$ and $\boldsymbol{\ \varepsilon}_{\mathrm{I},\mathrm{II}\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$. Using these for equation (4), we obtain
@@
$\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)+2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}$
$\hskip 10.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (14)
@@
$\approx-\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (15)
By means of equations (3), (5) and (8) one can express
$\cos\vartheta$ $\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (16)
$\sin\vartheta$ $\approx\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (17)
Now, using the relations between equations (14), (15) and (16), (17), the birefringent vector can finally be specified from equation (2):

### v3

@@ Abstract
Highly magnetized neutron stars have quantum refraction effects on pulsar emission due to the non-linearity of the quantum electrodynamics (QED) action. In this paper, we investigate the evolution of the polarization states of pulsar emission under the quantum refraction effects, combined with the dependence on the emission frequency, for dipole and quadrudipole pulsar models; we solve a system of evolution equations of the Stokes vector, where the birefringent vector, in which such effects are encoded, acts on the Stokes vector. At a fixed emission frequency, depending on the magnitude of the birefringent vector, dominated mostly by the magnetic field strength, the evolution of the Stokes vector largely exhibits three different patterns: (i) monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviors. These features are understood and confirmed by means of approximate analytical solutions to the evolution equations. Also, the evolution patterns are shown to differ between dipole and quadrudipole pulsar models, depending on the magnetic field strength.
@@
pulsars, magnetic fields, photon emission, photon polarization, nonlinear electrodynamics, nonlinear vacuum, vacuum birefringence, quantum refraction, polarization evolution
@@ 1 Introduction
Strong fields may open a window for testing fundamental physics. Even before quantum electrodynamics (QED) was fully developed and precisely tested in the weak field regime, Heisenberg and Euler showed that a strong electromagnetic field can polarize the Dirac vacuum Heisenberg1936Folgerungen . Schwinger introduced the proper-time integral method to obtain the one-loop effective QED action of the vacuum under a uniform electromagnetic field Schwinger1951Gauge . The so-called Heisenberg-Euler-Schwinger (HES) action provides an effective theory of electrodynamics in strong fields, in which the linear Maxwell vacuum is turned into a dielectric medium with electric, magnetic, and magneto-electric responses. Consequently, a photon propagating in a region of strong electromagnetic fields can experience vacuum birefringence, i.e., a quantum refraction effect Meszaros1992High ; Kim2023Vacuum . Furthermore, when the electric field is sufficiently strong to be comparable to the critical electric field strength ($E_{\rm{c}}=m_{\mathrm{e}}^{2}c^{3}/\left(e\hbar\right)\approx 1.3\times 10^{16}\,\rm{V/cm}$), electron-positron pairs can be created spontaneously out of the vacuum, which is called Sauter-Schwinger pair production Ruffini_2010 ; Fedotov:2022ely ; Hattori:2023egw . Observing these effects will validate the quantum vacuum model in the strong-field regime; however, vacuum birefringence is practically far more likely to be implemented or observed than Sauter-Schwinger pair production.
Although the relevant field strength is too high to attain by terrestrial means, an experiment and several proposals for testing strong-field QED have been reported. In the PVLAS (Polarizzazione del Vuoto con Laser, i.e., polarization of vacuum with laser) project, permanent, superconducting magnets have been used with laser as a probe, and a limit on vacuum birefringence has been reported for a field strength of $2.5\times 10^{4}\,\mathrm{G}$ Ejlli2020PVLASa . Recently, several proposals have appeared, in which ultra-intense laser fields are used with X-rays as a probe Karbstein2021Vacuum ; Shen2018Exploring ; Yu2023X ; the current ultra-intense laser can provide a magnetic field strength of $10^{10}\,\rm{G}$ Danson2019Petawatt , which corresponds to millisecond pulsars, but not to young pulsars. Although the fields from such lasers are still weaker than the critical magnetic field strength, $B_{\mathrm{c}}=m_{\mathrm{e}}^{2}c^{3}/e\hbar\approx 4.414\times 10^{13}\,\mathrm{G}$ by three orders, these proposals are promising for the observation of vacuum birefringence in the relatively weak-field regime. Thus, they are seriously considered to be conducted at upcoming ultra-intense laser facilities Danson2019Petawatt .
However, the observation of vacuum birefringence in the strong-field regime requires a field strength comparable to the critical value. It has been predicted that such extreme fields are available from astrophysical compact objects. For instance, highly magnetized neutron stars have magnetospheres whose field strength approaches up to $\sim 2\times 10^{15}\,\rm{G}$ (about 50 times as high as the critical field strength) Meszaros1992High ; Harding2006Physics . In this regard, several space telescope missions are being conducted or proposed to observe the X-rays from neutron stars for vacuum birefringence: the Imaging X-ray Polarimetry Explorer (IXPE) IXPE , X-ray Polarimeter Satellite (XPoSat) XPoSat , the enhanced X-ray Timing and Polarimetry (eXTP) Santangelo_2019 and the Compton Telescope project Wadiasingh2019Magnetars . The X-rays from a neutron star contain information about vacuum birefringence in its magnetosphere, and the birefringence effect accumulates over the magnetospheric size. Such accumulation is a great advantage compared to terrestrial laser experiments, not to mention the available field strength. This way, astrophysical compact objects can be used as a laboratory to test fundamental physics in the strong-field regime Raffelt1996Stars ; Meszaros1992High ; Kim2024Magnetars .
@@
The HES action is well approximated by the post-Maxwellian action, even up to the strength one order lower than the critical magnetic field $B_{\mathrm{c}}$, which keeps up to the quadratic terms of the Maxwell scalar and pseudo-scalar. Therefore, the post-Maxwellian action exhibits non-linear characteristics of vacuum polarization, such as quantum refraction Adler1971Photon ; Ni2013Foundations ; Denisov2016Pulsar . Previously, we have studied the quantum refraction effects on the propagation of a probe photon in the magnetic dipole field background of a pulsar model 10.1093/mnras/stae1304 . The study is non-trivial in comparison with other similar studies wherein the background magnetic field is assumed to be uniform, in that we have to deal with a dipole magnetic field, the strength and direction of which vary over space.
@@
In this work, we investigate the evolution of the polarization states of pulsar emission under the quantum refraction effects, combined with the dependence on the emission frequency, for dipole and quadrudipole (for the first time, to our knowledge) pulsar models; with growing theoretical and observational concerns for beyond-dipole effects, this study extends the scope of previous works by taking into account the multipolar magnetic field structure, the importance of which has been illuminated in different contexts of pulsar astronomy by a number of studies (see Gralla_2016 ; 10.1093/mnras/stz2524 ; Kazmierczak2019 ; Kalapotharakos_2021 , Petri2016Theory and references therein). To this end, we employ the evolution equations of the Stokes vector, where such effects are encoded into the birefringent vector that acts on the Stokes vector. The Stokes vector has a crucial advantage over the polarization vector in representing polarization states: it can be directly determined from experimentally measurable quantities and accommodate depolarization effects due to incomplete coherence and random processes during the photon propagation. Solutions of the evolution equations describe how the polarization states change along the photon propagation path from the emission point towards an observer. It turns out that the evolution of the Stokes vector, at a fixed frequency of emission, largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, dominated mostly by the magnetic field strength: (i) fractionally oscillatory - monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviors, which are found by numerical solutions and also confirmed by approximate analytical solutions. These are novel features rarely illuminated in previous studies on the same topic. In addition, it is investigated how the aforementioned features regarding the evolution of the Stokes vector change as we replace a dipole field with a quadrudipole field to modify the pulsar magnetic field structure. Throughout our analysis, X-ray emission from pulsars, with frequency $\sim 10^{18}\,\mathrm{Hz}$, is considered; in this regime, the vacuum contribution to the birefringence dominates that of the plasma Heyl2000 ; Wang2007Wave . Also, our analysis is sufficiently rigorous in solving the evolution equations of the Stokes vector, in that we feed into the equations the precise information of photon propagation under the pulsar rotation effect, through the magnetic field geometries of oblique dipole and quadrudipole rotators, with all the quantities involved fully affine-parameterized; then, the equations are solved solely in terms of an affine parameter.
@@
The paper is organized as follows. In Sect. 2.1, we introduce a system of evolution equations of the Stokes vector and apply this formalism to our pulsar emission model for an oblique dipole rotator. In Sect. 2.2, the evolution equations are solved for some known rotation-powered pulsars (RPPs) in three ways: fully numerically, via perturbation analysis, and using an analytical approximation. Also, we discuss the evolution patterns of the Stokes vectors resulting from the solutions. In Sect. 3.1, we consider a magnetic quadrudipole model for pulsar emission and look into the evolution equations under this model. In Sect. 3.2, we solve the evolution equations for the same RPPs fully numerically, and compare the results with those for the dipole case in Sect. 2.2. Then finally, we conclude the paper with discussions on other similar studies and future follow-up studies.
@@
2 Evolution of polarization states in strong magnetic field – dipole pulsars
@@ 2.1 Evolution equations of Stokes vector
Classically, polarization properties of pulsar emission are described by the Stokes parameters $\left\{I,Q,U,V\right\}$, where $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization of pulsar emission (for more details, see Appendix A). However, in the presence of a strong magnetic field in the background of the emission, the polarization evolves along the photon propagation path from the emission point towards an observer. The evolution of the polarization can be investigated systematically using the formalism initiated by Kubo1981 ; Kubo1983 ; Kubo1985 , and further developed by Heyl2000 ; Heyl2003highenergy ; Heyl2018Strongly ; Novak2018 , namely, a system of evolution equations of the Stokes vector, described as
@@
where $k\equiv\omega/c$ denotes the wave number for the electromagnetic radiation and $s$ is an affine parameter to measure the length of the photon trajectory, and $\mathbf{S}$ is the normalized Stokes vector, defined out of the Stokes parameters as $\mathbf{S}=\left(S_{1},S_{2},S_{3}\right)\allowbreak\equiv\left(Q/I,U/I,V/I\right)$, [^1] The classical Stokes vector can be expressed via pulse profiles of pulsar curvature emission, as illustrated in Appendix A. and $\mathbf{\hat{\Omega}}$ is the dimensionless birefringent vector, defined as [^2] Note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in the references above.
@@
with $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ being the two classical mode polarization vectors, orthogonal to each other and to $\mathbf{\hat{n}}_{\left[0\right]}$; the specific forms of $\mathbf{\hat{n}}_{\left[0\right]}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ are later given by Eqs. (8), (12) and (13), respectively, for the magnetic field of an oblique dipole rotator as described by Eq. (5) and Fig. 1.
Figure 1: A cross-sectional view of a pulsar magnetosphere with the dipole magnetic field lines (green) around a neutron star. The vertical dashed line (black) and the inclined solid line (red) represent the rotation axis and the magnetic axis, respectively. $\alpha$ between these axes denotes the inclination angle. The scale of the unity in this graph is equivalent to the neutron star radius $\sim 10^{6}\,\mathrm{cm}$. The red dashed line represents the trajectory curve of the light ray traced by the propagation vector $\mathbf{\hat{n}}_{\left[0\right]}$ as projected onto the $xz$-plane. (Credit: Kim2021Generala , reproduced with modifications.)
@@
$\mathbf{B}\left(r,\theta,\phi\right)$ $=\frac{2\mu\left(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{r}}+\frac{\mu\left(\cos\alpha\sin\theta-\sin\alpha\cos\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{\theta}}$
$\hskip 12.0pt+\frac{\mu\sin\alpha\sin\phi\,}{r^{3}}\mathbf{e}_{\hat{\phi}},$ (5)
@@
However, at the same time, our pulsar magnetosphere rotates, and therefore the field lines get twisted due to the magneto-centrifugal acceleration on the plasma particles moving along the field lines Blandford1982 . Then, taking into consideration this magneto-hydrodynamic (MHD) effect, the direction of the classical photon propagation, which must line up with the particle velocity in order for an observer to receive the radiation, can be described as Gangadhara2005
$\mathbf{\hat{n}}_{\left[0\right]}=\beta\mathbf{\hat{B}}+\frac{\mathbf{\Omega}_{\mathrm{p}}\times\mathbf{r}}{c},$ (6)
@@
During the rotation the azimuthal phase changes by $\phi\sim\Omega_{\mathrm{p}}t$, while our photon has propagated a distance by $s\sim ct$. In our analysis, the photon propagation is described with the consideration of the MHD effect above, assuming $\phi$ to be very small; e.g., $\phi\lesssim 10^{-1}$ is considered for a millisecond pulsar with $\Omega_{\mathrm{p}}\sim 10^{2}\,\mathrm{Hz}$, during the time of rotation $t\lesssim 10^{-3}\,\mathrm{s}$, such that $s\lesssim 10^{7}\,\mathrm{cm}$, which corresponds to the propagation distance within about $10$ times the neutron star radius. For Eq. (6) we take only the leading order expansions of $\mathbf{\hat{B}}\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ and $\beta\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ in $\phi$ from Eqs. (5) and (7), respectively, and can express the classical propagation vector $\mathbf{\hat{n}}_{\left[0\right]}$ in Cartesian coordinates as
@@
where we have considered $\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\lesssim\phi$, e.g., for a millisecond pulsar with $\Omega_{\mathrm{p}}\sim 10^{2}\,\mathrm{Hz}$ and $r_{\mathrm{o}}\sim 10^{6}\,\mathrm{cm}$, such that $\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\lesssim\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\lesssim\phi^{2}$, all to be ignored in our analysis, and have substituted $\phi=\Omega_{\mathrm{p}}s/c$ in Eq. (11), the leading order rotational effect to be considered in our analysis.
The orthogonal pair of classical mode polarization vectors, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$, both being also orthogonal to $\mathbf{\hat{n}}_{\left[0\right]}$ as given by Eq. (8) above, are determined as
@@
such that the three vectors, $\mathbf{\hat{n}}_{\left[0\right]}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ form an orthonormal basis. [^5] It can be checked out that $\mathbf{\hat{n}}_{\left[0\right]}\cdot\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}\approx 0+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$, $\mathbf{\hat{n}}_{\left[0\right]}\cdot\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}=0$ and $\boldsymbol{\varepsilon}_{\,\mathrm{I}\left[0\right]}\cdot\boldsymbol{\ \varepsilon}_{\mathrm{II}\left[0\right]}=0$ while $\mathbf{\hat{n}}_{\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$ and $\boldsymbol{\ \varepsilon}_{\mathrm{I},\mathrm{II}\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$. Using these for Eq. (4), we obtain
@@
$\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)+2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\hat{n}_{y\left[0\right]}$
$\hskip 12.0pt-\frac{\Omega_{\mathrm{p}}\sin\alpha\,s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (14)
@@
$\approx-\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}$
$\hskip 12.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (15)
By means of Eqs. (3), (5) and (8) one can express
$\cos\vartheta$ $\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}$
$\hskip 12.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (16)
$\sin\vartheta$ $\approx\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}$
$\hskip 12.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (17)
Now, using the relations between Eqs. (14)-(17), the birefringent vector can finally be specified from Eq. (2):